### v2

# Scaling Population-Based Reinforcement Learning with GPU Accelerated Simulation

Abstract
In recent years, deep reinforcement learning (RL) has shown its effectiveness in solving complex continuous control tasks like locomotion and dexterous manipulation. However, this comes at the cost of an enormous amount of experience required for training, exacerbated by the sensitivity of learning efficiency and the policy performance to hyperparameter selection, which often requires numerous trials of time-consuming experiments. This work introduces a Population-Based Reinforcement Learning (PBRL) approach that exploits a GPU-accelerated physics simulator to enhance the exploration capabilities of RL by concurrently training multiple policies in parallel. The PBRL framework is applied to three state-of-the-art RL algorithms – PPO, SAC, and DDPG – dynamically adjusting hyperparameters based on the performance of learning agents. The experiments are performed on four challenging tasks in Isaac Gym – Anymal Terrain, Shadow Hand, Humanoid, Franka Nut Pick – by analyzing the effect of population size and mutation mechanisms for hyperparameters. The results show that PBRL agents achieve superior performance, in terms of cumulative reward, compared to non-evolutionary baseline agents. The trained agents are finally deployed in the real world for a Franka Nut Pick task, demonstrating successful sim-to-real transfer. Code and videos of the learned policies are available on our project website.
Index Terms:
Reinforcement Learning, Dexterous Manipulation, Evolutionary Learning, Sim-to-RealI Introduction
Many domains have seen tremendous advancements of reinforcement learning (RL) applications in recent years, ranging from playing challenging games [1, 2] to learning high-dimensional continuous control in robotics [3, 4, 5]. Tasks such as dexterous manipulation [6], legged locomotion [7], and mobile navigation [8] have been learned using deep RL. A primary challenge in training RL policies is the need for large amounts of training data. RL methods rely on effective exploration to discover control policies, which can be particularly challenging when operating in high-dimensional continuous spaces [9]. Moreover, the performance of the learned policy is highly dependent on the tedious tuning of hyperparameters. Hyperparameter tuning, in fact, can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment. One way to deal with the problem of data inefficiency is to train in simulation before transferring to reality [10, 11, 12]. However, the time required to train the policy in simulation increases significantly with the task complexity. For example, in [11], learning a block re-orientation task with a robot hand took around 14 days and enormous computing resources. In addition, policies trained in simulation often fail to perform on a real system due to discrepancies between the simulation and the real world. Recent advances in GPU-accelerated simulation, such as Isaac Gym [13, 14], have made it possible to run thousands of parallel environments on a single GPU, which reduces the training times significantly. However, successfully training RL policies still requires carefully tuned hyperparameters to explore efficiently.
I-A Related Works
I-A1 Massively Parallel Simulation
The advent of GPU-based simulation has significantly improved simulation throughput by incorporating massive parallelism on a single GPU [13, 15]. A number of recent works have exploited this parallelism to demonstrate impressive performance on challenging control problems using RL [14, 16, 17]. However, almost all recent works use the same algorithm, i.e. Proximal Policy Optimization (PPO) [18] to train RL policies; other common approaches include off-policy techniques, e.g. Soft Actor-Critic (SAC) [19] and Deep Deterministic Policy Gradient (DDPG) [20]. While simple and effective, all these algorithms require a range of hyperparameters that need to be tuned for each task to ensure sufficient exploration and stabilize training.
I-A2 Population-Based Reinforcement Learning
Population-based approaches offer a promising solution to deal with exploration and hyperparameter tuning by training a set of policies as opposed to a single policy. Multiple agents can be used to collect diverse experiences that improve robustness and stabilize training by dynamically adapting the hyperparameters. Some prior works have shown remarkable results in employing these approaches to train deep RL policies in domains like strategy games and multi-agent interaction [21, 22, 23]. However, there is almost no existing research investigating PBRL methods for robotics. This is due to the fact that the computational complexity and training time of these approaches increase linearly with the number of agents on CPU-based simulators like MuJoCo [24], requiring multiple worker machines with separate simulation instances to speed up data collection. Isaac Gym allows simulating thousands of robots in parallel, giving access to a vast amount of experience data, rendering it suitable to efficiently train a population of RL agents.
Training various RL agents provides a mechanism for meta-optimization, utilizing the potential of both learning and evolution [25]. One successful example of PBRL methods is population-based training (PBT) [26], which allows training multiple policies concurrently to enhance the exploration capabilities of the agents in generating diverse behaviors. In PBT, the faster inner loop of policy optimization is ultimately evaluated by the slower-running outer loop of evolutionary selection. PBT trains a population of agents with different hyperparameters and uses a genetic algorithm to update the population periodically. Recently, DexPBT, a decentralized PBRL approach has been proposed to learn dynamic manipulation between two hand-arm systems using parallel simulations [27]. The authors developed a decentralized implementation to evolve agents in distributed computing environments using on-policy RL, achieving impressive results in dexterous manipulation. However, sim-to-real transfer has not been performed, highlighting the complexity of deploying policies on real systems.
In contrast, this work targets a broader range of real-world tasks including locomotion and manipulation, and transfers the policy onto a real robot without any adaptation phase. In addition, the PBRL framework is successfully applied to both off-policy (SAC, DDPG) and on-policy (PPO) RL algorithms, analyzing the implications of critical design choices, i.e., the number of agents and the mutation mechanisms.
I-A3 Sim-to-Real Transfer
Despite the calibration efforts to model the physical system accurately, simulation is still a rough approximation. The differences between the dynamics of simulated and real systems cause a “reality gap” that makes it unlikely for a simulation-trained policy to successfully transfer to a physical system.
In literature, researchers have put a significant effort into diminishing this gap: to this aim, most of the approaches leverage domain randomization [4, 6, 16, 17, 28, 29] to expose the policy to a wide range of observation distributions in simulation, thus improving generalization onto a real system. Nevertheless, naive domain randomization might not be sufficient to completely attenuate the dynamics gap: for instance, [30] employs a specific network to mimic the real actuation system. Another technique in this context is policy-level action integrator (PLAI) [10], a simple yet effective algorithm aimed at compensating the sim-to-real dynamic discrepancies with an integral action, which proved to be paramount for a successful transfer.
In this paper, we employ sim-to-real strategies to deploy a policy on a real system; to the best of the authors’ knowledge, this work represents the first instance of deploying PBRL agents on real hardware.
I-B Contribution
This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation in order to tune the hyperparameters during training. The PBRL framework is evaluated on four complex tasks in locomotion and manipulation available in Isaac Gym [13]: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1). The results show that better performance is achieved when training a population of agents compared to a single-agent baseline on all tasks. The comparison is provided across 3 RL algorithms (PPO, SAC, and DDPG), varying the number of agents in a population, and across different hyperparameter mutation schemes. Finally, the PBRL agents are deployed on a real Franka Panda robot for a Franka Nut Pick task, without any policy adaptation phase on the physical system. In summary, the main contributions of this work are:
a population-based RL framework that utilizes GPU-accelerated simulation to train robotic manipulation tasks by adaptively optimizing the set of hyperparameters during training;
simulations demonstrating the effectiveness of the PBRL approach on 4 tasks using 3 RL algorithms, including both on-policy and off-policy methods, investigating the performance w.r.t. the number of agents and mutation mechanisms;
sim-to-real transfer of PBRL policies onto a real Franka Panda robot;
an open-source codebase to train policies using the PBRL algorithm.
II Methods
This section describes the core concepts involved in the PBRL framework. The overall approach, illustrated in Figure 2, can be viewed as a multi-layered training process consisting of an inner optimization loop with RL and an outer loop of online evolutionary selection with population-based training. During training, the parameters of the agent’s policy are updated at a higher rate using RL than the hyperparameters defining the RL procedure.
II-A Reinforcement Learning
The RL problem is modeled as a Markov Decision Process (MDP), where an agent interacts with the environment in order to maximize the expected sum of episodic rewards. Specifically, an MDP is defined as , where is the set of states, is the set of actions, is the transition dynamics, i.e., , where defines the set of a probability distribution over , is the reward function, and represents the discount factor. The goal is formulated as learning a policy, either stochastic, , or deterministic, , where represents a probability distribution over and encapsulates the policy parameters, whose cardinality depends on the selected algorithm and network architecture. In this work, the policy is learned using the on-policy method PPO, or either of the off-policy methods SAC or DDPG. All these algorithms use an actor-critic architecture simultaneously learning the policy (actor) and the value function approximators (critics) . The implementation of critics in SAC and DDPG relies on double Q-learning and -step returns.
To train the policy with PPO, a learning rate (LR) adaptation procedure is used based on a Kullback–Leibler (KL) divergence starting from an initial value [13]. At the end of each update iteration, the LR is increased by a factor of when the KL divergence between the current policy and the old policy is below the specified threshold, or reduced by if the KL divergence exceeds the threshold.
In DDPG, the common practice involves adding a small noise to the deterministic actions of the policy to enable exploration. In this work, the noise is added following a mixed exploration strategy [31], where the general idea is akin to adding a different noise level for each environment when training in a massively parallel regime. For the -th environment out of environments, a zero mean and uncorrelated Gaussian noise is given as: , where .
Hyperparameter | PPO | PBRL-PPO | |||||||
Anymal Terrain |
Franka Nut Pick | Anymal Terrain |
Franka Nut Pick | ||||
Environments per agent | 4096 | 16384 | 128 | 1024 | 4096 | 128 | |||
MLP hidden units | [512, 256, 128] | [512, 256, 128] | [256, 128, 64] | [512, 256, 128] | [512, 256, 128] | [256, 128, 64] | |||
Horizon | 32 | 16 | 120 | 32 | 16 | 120 | |||
Batch size | 8192 | 32768 | 512 | 8192 | 8192 | 512 | |||
Actor variance | 0.5 | 1 | 1 | 0.3 – 1 | 0.3 – 1 | 0.3 – 1 | |||
KL threshold | 0.016 | 0.016 | 0.016 | 0.08 – 0.016 | 0.08 – 0.016 | 0.08 – 0.016 | |||
Entropy loss coefficient | 0.001 | 0.001 | 0 | 0.0001 – 0.001 | 0.0001 – 0.001 | 0.0001 – 0.001 | |||
Epochs | 8 | 4 | 8 | 8 | 4 | 8 | |||
Discount factor | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | |||
GAE lambda | 0.95 | 0.95 | 0.95 | 0.95 | 0.95 | 0.95 | |||
PPO clip | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | 0.2 | |||
Initial LR | |||||||||
LR adaptation gain | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 |
II-B Population-Based Training
In standard RL, the agent aims to learn an optimal policy by interacting with an environment and iteratively updating the policy through some kind of optimization method. In contrast, PBRL uses a population of agents , each interacting with the environment independently to collect experience and learn its own policy. Using evolutionary selection, the population is evaluated periodically based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
In this work, a specific PBRL approach, population-based training (PBT), is employed as an outer optimization loop to enable diverse exploration and dynamically adapt the hyperparameters in high-dimensional continuous control tasks. Each agent is characterized by a vector and the set of hyperparameters , where contains the parameters of the policy, and contains the hyperparameters that are optimized during training. To represent the whole population , we denote with , and the sets of all the parameters, hyperparameters and policies respectively.
Hyperparameter |
Environments per agent* | 2048 | 2048 | ||||
MLP hidden units* | [512, 256, 128] | [512, 256, 128] | ||||
Batch size* | 4096 | 4096 | ||||
Horizon | 1 | 1 | ||||
Target update rate | ||||||
Actor learning rate | 0.0001 | 0.0001 – 0.001 | ||||
Critic learning rate | 0.0001 | 0.0001 – 0.001 | ||||
DDPG exploration | 0.01 | 0.01 – 0.1 | ||||
DDPG exploration | 1 | 0.5 – 1 | ||||
SAC target entropy | -20 | -20 – -10 | ||||
Replay buffer size | ||||||
Epochs | 4 | 4 | ||||
-step returns | 3 | 3 |
Algorithm 1 provides pseudocode for the PBRL. The training process runs in iterations, where all agents are first independently trained by performing updates to the vector . After a certain number of policy updates (each agent having been trained for some steps), the agents are evaluated and sorted based on the average return obtained over all of the previous episodes. The agents in get replaced by randomly-sampled agents in with mutated hyperparameters, while the rest of the agents in and continue training.
To generate the mutated hyperparameters, 3 mutation mechanisms are considered (see line 15 of Algorithm 1): (i) random perturbation is applied to the hyperparameters of the parent agent(s) through perturbation factors in Table III; (ii) new hyperparameters are sampled from a prior uniform distribution with bounds specified in Table I and II; (iii) according to the DexPBT mutation scheme [27], hyperparameters are multiplied or divided by a random number sampled from a uniform distribution, i.e., with probability . Section III-B4 compares all 3 mutation schemes. After beginning the training, evolution is enabled after steps as in [22] to allow for initial exploration and promote population diversity.
III Experiments
III-A Environments
The PBRL framework is evaluated on some of the most challenging benchmark tasks available in Isaac Gym, including Anymal Terrain, Shadow Hand, Humanoid and Franka Nut Pick (Figure 1). The experiments are conducted on a workstation with a single NVIDIA RTX 4090 GPU and of RAM. Parallelizing the data collection across the GPU, Isaac Gym’s PhysX engine can simulate thousands of environments using the above hardware.
III-B Results
Parameter | Value | |
Franka Nut Pick | Others | |
Evolution start | steps | steps |
Evolution frequency | steps | steps |
Perturbation factor (min.) | 0.8 | 0.8 |
Perturbation factor (max.) | 1.2 | 1.2 |
PPO PBRL-PPO 4 agents
PBRL-PPO 8 agents PBRL-PPO 16 agents
The experiments focus on optimizing the hyperparameters of the RL agents in a population and comparing the results against non-evolutionary baseline agents. For each case of baseline agents, 4 experiments are run with different seeds. Table I and II provide the hyperparameters for on-policy and off-policy algorithms, listing the sampling ranges of those optimized through the PBRL Algorithm 1. The initial values for each agent are sampled uniformly from a prior distribution with a given range.
III-B1 PBRL-PPO
For the PPO agents, the tuned hyperparameters are the KL divergence threshold for an adaptive LR, the entropy loss coefficient, and the variance of action selection. These parameters are crucial in ensuring sufficient exploration of the environment. Figure 3 shows the learning curves for the single-agent PPO baseline and PBRL-PPO for . The results demonstrate that PBRL-PPO outperforms PPO on 3 out of 4 tasks, yielding a higher return, with significant improvement seen in Anymal Terrain, which involves traversing increasingly challenging terrain. For Franka Nut Pick, PBRL agents achieve comparable performance to the baseline PPO agents. This is because, in this relatively straightforward task, randomization alone suffices for a thorough exploration of the state/action space.
III-B2 PBRL-SAC
The hyperparameters optimized in PBRL-SAC include the LR of the actor-critic networks and the target entropy factor to control the exploration behavior. Entropy is a central theme in SAC agents as the policy is trained to maximize the trade-off between the expected return and the entropy. The experiments in PBRL-SAC are run with a population size of . The replay buffers in off-policy methods require more memory allocation on GPU than on-policy methods. Since all agents in PBRL-SAC train with their individual replay buffers, the maximum population size is set to 8 due to the limited GPU memory. The training performance of SAC and PBRL-SAC is shown in Figure 4. PBRL-SAC improves the training performance compared to non-evolutionary SAC on 3 out of 4 tasks, yielding a remarkable improvement on both Shadow Hand and Franka Nut Pick, while comparable results are achieved on Humanoid, probably due to the limited task complexity.
III-B3 PBRL-DDPG
SAC PBRL-SAC 4 agents PBRL-SAC 8 agents
DDPG PBRL-DDPG 4 agents PBRL-DDPG 8 agents
In DDPG, exploration noise is added to the output of the deterministic actor. As mentioned in Section II-A, different noise levels are added for different environments uniformly within the range . Both these parameters are crucial in controlling the amount of exploration in DDPG agents. In PBRL-DDPG, the hyperparameters optimized during training include the minimum and the maximum bounds for noise levels, i.e., , and the LRs of the actor and the critic. As in PBRL-SAC, the maximum population size in PBRL-DDPG is set to 8 due to the presence of independent replay buffers and GPU memory limitations. Figure 5 shows that PBRL-DDPG achieves significantly better training performance than DDPG on all 4 benchmark tasks.
III-B4 Mutation Comparison
Figure 6 shows the results using 3 different mutation schemes for PBRL-PPO and PBRL-DDPG. As mentioned in Section II-B, the hyperparameters for under-performing agents are generated either by sampling from an original prior distribution, by perturbing the parent’s values through perturbation factors given in Table III, or through the DexPBT mutation scheme presented in [27]. In the latter, the hyperparameters have a probability of getting multiplied or divided by a random number sampled from the uniform distribution, . The results show that the perturbed agents either exceed or are on par with the performance of other mutation schemes in 6 out of 8 experiments. The DexPBT mutation scheme performs better with PBRL-DDPG on Humanoid and Franka Nut Pick tasks, which are less challenging compared to others. The combination of two mutation schemes might discover better exploration strategies for a wider range of tasks. Analyzing the potential synergies between the two remains a prospect for future investigation.
Parameter | Value |
Franka initial position | |
Franka initial rotation | |
Nut initial position | [0.42, 0.27, 0.02] [0.18, 0.13, 0.01] |
TSI proportional gains | [1000, 1000, 1000, 50, 50, 50] |
TSI derivative gains | [63.25, 63.25, 63.25, 1.414, 1.414, 1.414] |
Action scale | 0.0001 |
III-C Sim-to-Real Transfer
In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation on a physical robot, deploying the actions with PLAI [10]. The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
Compared to the original task [4], we applied the following changes to make the simulated environment as close as possible to the real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to resemble the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose. The parameters used in the simulated environment and the real controller are reported in Table IV.
During experiments, the following policies were deployed, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO. With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase. As shown in Figure 7, PBRL-PPO with has the highest success rate. Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs. ), also highlighting lower variance compared to both baseline PPO and PBRL-PPO with .
Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V. Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation. This demonstrates that PBRL-PPO, favoring exploration across the agents, yields a successful real-world transfer, despite reaching similar rewards as the standard PPO. Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
PPO PBRL-PPO 4 agents
PBRL-PPO 8 agents PBRL-PPO 16 agents
Algorithm | Agent | Successful trials | Success rate |
PBRL-PPO | Best | 27/30 | 90% |
PBRL-PPO | Worst | 21/30 | 70% |
PPO | Best | 19/30 | 63.33% |
III-D Discussion
While the PBRL agents perform better than the non-evolutionary agents in almost all the experiments, the impact of population size across RL algorithms and tasks shows no consistent pattern. One may hypothesize that larger and more diverse populations might lead to a better final performance. However, the results in this work indicate that using a larger population size does not necessarily yield substantial benefits for every task. This is in contrast to the common belief that population-based methods rely on larger population sizes to effectively explore the hyperparameter space [26, 35]. The optimal population size, instead, depends on various factors, including task complexity, RL algorithm, and interaction dynamics among agents. While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice. Perhaps, larger population sizes also perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
Additionally, the performance of PBRL may be lower than non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori. This can be noticed on a Humanoid task trained with SAC in Figure 4: indeed, baseline policies achieve a higher reward than PBRL-SAC with 4 agents; nevertheless, 8 agents are capable of outperforming the baseline. Thus, the benefits provided by PBRL will become more apparent for new tasks where ideal hyperparameter ranges are not known in advance. In this sense, PBRL can be thought of as an exploratory approach to search for unknown optimal configurations of newly designed tasks.
IV Conclusion
IV-A Summary
In this paper, a PBRL framework has been employed for training a population of RL agents by making use of high-throughput GPU-accelerated simulation. The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes. The results showed the effectiveness of PBRL in improving final performance by adapting the hyperparameters online. PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning. Finally, we release the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms on challenging robotic tasks.
IV-B Future Works
Many interesting directions remain for future research. An immediate extension could be to use a dedicated fitness metric for each population sub-group to prioritize long-term performance [36]. This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training.
Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation, and assembly) in the real world with sim-to-real techniques. With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable, accounting for diverse skills labeling and learning.

### v3

@@
 The experiments are performed on four challenging tasks in Isaac Gym – Anymal Terrain, Shadow Hand, Humanoid, Franka Nut Pick – by analyzing the effect of population size and mutation mechanisms for hyperparameters.
-The results show that PBRL agents achieve superior performance, in terms of cumulative reward, compared to non-evolutionary baseline agents.
-The trained agents are finally deployed in the real world for a Franka Nut Pick task, demonstrating successful sim-to-real transfer.
+The results demonstrate that PBRL agents outperform non-evolutionary baseline agents across tasks essential for humanoid robots, such as bipedal locomotion, manipulation, and grasping in unstructured environments.
+The trained agents are finally deployed in the real world for the Franka Nut Pick manipulation task.
+To our knowledge, this is the first sim-to-real attempt for successfully deploying PBRL agents on real hardware.
 Code and videos of the learned policies are available on our project website.
-Index Terms:
-Reinforcement Learning, Dexterous Manipulation, Evolutionary Learning, Sim-to-RealI Introduction
+I Introduction
 Many domains have seen tremendous advancements of reinforcement learning (RL) applications in recent years, ranging from playing challenging games [1, 2] to learning high-dimensional continuous control in robotics [3, 4, 5].
@@
 Moreover, the performance of the learned policy is highly dependent on the tedious tuning of hyperparameters.
-Hyperparameter tuning, in fact, can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment.
+Hyperparameter tuning can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment.
 One way to deal with the problem of data inefficiency is to train in simulation before transferring to reality [10, 11, 12].
@@
 One successful example of PBRL methods is population-based training (PBT) [26], which allows training multiple policies concurrently to enhance the exploration capabilities of the agents in generating diverse behaviors.
-In PBT, the faster inner loop of policy optimization is ultimately evaluated by the slower-running outer loop of evolutionary selection.
 PBT trains a population of agents with different hyperparameters and uses a genetic algorithm to update the population periodically.
@@
 I-B Contribution
-This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation in order to tune the hyperparameters during training.
-The PBRL framework is evaluated on four complex tasks in locomotion and manipulation available in Isaac Gym [13]: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1).
+This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation to dynamically adjust the hyperparameters during training.
+We evaluate the PBRL framework on four complex tasks that require learning essential skills for humanoid robots: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1), available in Isaac Gym [13].
 The results show that better performance is achieved when training a population of agents compared to a single-agent baseline on all tasks.
@@
 | Initial LR | |||||||||
-LR adaptation gain | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 |
+| LR adaptation gain | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 | 1.5 | |||
 II-B Population-Based Training
@@
 In contrast, PBRL uses a population of agents , each interacting with the environment independently to collect experience and learn its own policy.
-Using evolutionary selection, the population is evaluated periodically based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
+Using evolutionary selection, the population is periodically evaluated based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
 In this work, a specific PBRL approach, population-based training (PBT), is employed as an outer optimization loop to enable diverse exploration and dynamically adapt the hyperparameters in high-dimensional continuous control tasks.
@@
 | Epochs | 4 | 4 | ||||
--step returns | 3 | 3 |
+| -step returns | 3 | 3 |
 Algorithm 1 provides pseudocode for the PBRL.
@@
 For each case of baseline agents, 4 experiments are run with different seeds.
-Table I and II provide the hyperparameters for on-policy and off-policy algorithms, listing the sampling ranges of those optimized through the PBRL Algorithm 1.
-The initial values for each agent are sampled uniformly from a prior distribution with a given range.
+Tables I and II list the hyperparameters for on-policy and off-policy algorithms, including the sampling ranges of those optimized through the PBRL Algorithm 1.
+The initial values for each agent are uniformly sampled from a prior distribution with a given range.
 III-B1 PBRL-PPO
@@
 This is because, in this relatively straightforward task, randomization alone suffices for a thorough exploration of the state/action space.
+SAC PBRL-SAC 4 agents PBRL-SAC 8 agents
+DDPG PBRL-DDPG 4 agents PBRL-DDPG 8 agents
 III-B2 PBRL-SAC
-The hyperparameters optimized in PBRL-SAC include the LR of the actor-critic networks and the target entropy factor to control the exploration behavior.
-Entropy is a central theme in SAC agents as the policy is trained to maximize the trade-off between the expected return and the entropy.
-The experiments in PBRL-SAC are run with a population size of .
-The replay buffers in off-policy methods require more memory allocation on GPU than on-policy methods.
-Since all agents in PBRL-SAC train with their individual replay buffers, the maximum population size is set to 8 due to the limited GPU memory.
+In PBRL-SAC, the optimized hyperparameters include the LR of the actor-critic networks and the target entropy factor.
+Entropy is key in SAC agents as the policy is trained to maximize the trade-off between the expected return and exploration.
+Experiments are run with a population size of .
+Due to higher memory needs for replay buffers in off-policy methods, the maximum population size is limited to 8.
 The training performance of SAC and PBRL-SAC is shown in Figure 4.
@@
 III-B3 PBRL-DDPG
-SAC PBRL-SAC 4 agents PBRL-SAC 8 agents
-DDPG PBRL-DDPG 4 agents PBRL-DDPG 8 agents
 In DDPG, exploration noise is added to the output of the deterministic actor.
@@
 III-C Sim-to-Real Transfer
-In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation on a physical robot, deploying the actions with PLAI [10].
+In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation, executing the actions with PLAI [10].
 The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
-Compared to the original task [4], we applied the following changes to make the simulated environment as close as possible to the real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to resemble the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose.
+Compared to the original task [4], we applied the following changes to make the simulated environment resemble real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to reflect the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose.
 The parameters used in the simulated environment and the real controller are reported in Table IV.
@@
 Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs.
-), also highlighting lower variance compared to both baseline PPO and PBRL-PPO with .
+).
 Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V.
 Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation.
-This demonstrates that PBRL-PPO, favoring exploration across the agents, yields a successful real-world transfer, despite reaching similar rewards as the standard PPO.
+Unlike the baseline PPO agent, which continued to produce small movements after reaching the target, PBRL-PPO agents remained more stable, leading to a higher success rate.
+This demonstrates that PBRL agents, while achieving similar rewards to a single agent, learn behaviors that exhibit greater robustness to environment variability due to the diversity in agent populations.
 Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
@@
 While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice.
-Perhaps, larger population sizes also perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
+Larger population sizes seem to perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
 Additionally, the performance of PBRL may be lower than non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori.
@@
 IV-A Summary
-In this paper, a PBRL framework has been employed for training a population of RL agents by making use of high-throughput GPU-accelerated simulation.
-The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes.
-The results showed the effectiveness of PBRL in improving final performance by adapting the hyperparameters online.
+In this paper, a PBRL framework has been employed to train a population of RL agents by making use of high-throughput GPU-accelerated simulation.
+The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of locomotion and manipulation benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes.
+The results showed the effectiveness of PBRL in improving final performance through online adaptation of hyperparameters.
 PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning.
-Finally, we release the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms on challenging robotic tasks.
+Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms.
 IV-B Future Works
@@
 This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training.
-Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation, and assembly) in the real world with sim-to-real techniques.
-With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable, accounting for diverse skills labeling and learning.
+Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation or assembly) in the real world with sim-to-real techniques.
+With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable.

### v4

@@
-Scaling Population-Based Reinforcement Learning with GPU Accelerated Simulation
+Benchmarking Population-Based Reinforcement Learning across Robotic Tasks with GPU-Accelerated Simulation
 Abstract
-In recent years, deep reinforcement learning (RL) has shown its effectiveness in solving complex continuous control tasks like locomotion and dexterous manipulation.
+In recent years, deep reinforcement learning (RL) has shown its effectiveness in solving complex continuous control tasks.
 However, this comes at the cost of an enormous amount of experience required for training, exacerbated by the sensitivity of learning efficiency and the policy performance to hyperparameter selection, which often requires numerous trials of time-consuming experiments.
-This work introduces a Population-Based Reinforcement Learning (PBRL) approach that exploits a GPU-accelerated physics simulator to enhance the exploration capabilities of RL by concurrently training multiple policies in parallel.
-The PBRL framework is applied to three state-of-the-art RL algorithms – PPO, SAC, and DDPG – dynamically adjusting hyperparameters based on the performance of learning agents.
+This work leverages a Population-Based Reinforcement Learning (PBRL) approach and a GPU-accelerated physics simulator to enhance the exploration capabilities of RL by concurrently training multiple policies in parallel.
+The PBRL framework is benchmarked against three state-of-the-art RL algorithms – PPO, SAC, and DDPG – dynamically adjusting hyperparameters based on the performance of learning agents.
 The experiments are performed on four challenging tasks in Isaac Gym – Anymal Terrain, Shadow Hand, Humanoid, Franka Nut Pick – by analyzing the effect of population size and mutation mechanisms for hyperparameters.
-The results demonstrate that PBRL agents outperform non-evolutionary baseline agents across tasks essential for humanoid robots, such as bipedal locomotion, manipulation, and grasping in unstructured environments.
-The trained agents are finally deployed in the real world for the Franka Nut Pick manipulation task.
-To our knowledge, this is the first sim-to-real attempt for successfully deploying PBRL agents on real hardware.
+The results show that PBRL agents achieve superior performance, in terms of cumulative reward, compared to non-evolutionary baseline agents.
+Moreover, the trained agents are finally deployed in the real world for a Franka Nut Pick task.
+To our knowledge, this is the first sim-to-real attempt for deploying PBRL agents on real hardware.
 Code and videos of the learned policies are available on our project website.
-I Introduction
+Index Terms:
+Reinforcement Learning, Evolutionary Learning, Sim-to-Real, Dexterous ManipulationI Introduction
 Many domains have seen tremendous advancements of reinforcement learning (RL) applications in recent years, ranging from playing challenging games [1, 2] to learning high-dimensional continuous control in robotics [3, 4, 5].
@@
 RL methods rely on effective exploration to discover control policies, which can be particularly challenging when operating in high-dimensional continuous spaces [9].
-Moreover, the performance of the learned policy is highly dependent on the tedious tuning of hyperparameters.
-Hyperparameter tuning can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment.
+Moreover, the performance of the learned policy is highly dependent on the tedious hyperparameters tuning procedure, a very time-consuming process often requiring several manual trials to determine the best values for the specific task and the learning environment.
 One way to deal with the problem of data inefficiency is to train in simulation before transferring to reality [10, 11, 12].
@@
 In addition, policies trained in simulation often fail to perform on a real system due to discrepancies between the simulation and the real world.
-Recent advances in GPU-accelerated simulation, such as Isaac Gym [13, 14], have made it possible to run thousands of parallel environments on a single GPU, which reduces the training times significantly.
-However, successfully training RL policies still requires carefully tuned hyperparameters to explore efficiently.
+Recent advances in GPU-accelerated simulation, such as Isaac Gym [13, 14], have made it possible to run thousands of parallel environments on a single GPU, which significantly reduces the training times.
+Nevertheless, successfully training RL policies still requires carefully tuned hyperparameters to explore efficiently.
 I-A Related Works
-I-A1 Massively Parallel Simulation
-The advent of GPU-based simulation has significantly improved simulation throughput by incorporating massive parallelism on a single GPU [13, 15].
+I-A1 Massively-Parallel Simulation
+The advent of GPU-based simulation has significantly improved simulation throughput by enabling massive parallelism on a single GPU [13, 15].
 A number of recent works have exploited this parallelism to demonstrate impressive performance on challenging control problems using RL [14, 16, 17].
-However, almost all recent works use the same algorithm, i.e.
+Almost all recent works use the same algorithm, i.e.
 Proximal Policy Optimization (PPO) [18] to train RL policies; other common approaches include off-policy techniques, e.g.
@@
 While simple and effective, all these algorithms require a range of hyperparameters that need to be tuned for each task to ensure sufficient exploration and stabilize training.
-I-A2 Population-Based Reinforcement Learning
-Population-based approaches offer a promising solution to deal with exploration and hyperparameter tuning by training a set of policies as opposed to a single policy.
-Multiple agents can be used to collect diverse experiences that improve robustness and stabilize training by dynamically adapting the hyperparameters.
-Some prior works have shown remarkable results in employing these approaches to train deep RL policies in domains like strategy games and multi-agent interaction [21, 22, 23].
-However, there is almost no existing research investigating PBRL methods for robotics.
-This is due to the fact that the computational complexity and training time of these approaches increase linearly with the number of agents on CPU-based simulators like MuJoCo [24], requiring multiple worker machines with separate simulation instances to speed up data collection.
-Isaac Gym allows simulating thousands of robots in parallel, giving access to a vast amount of experience data, rendering it suitable to efficiently train a population of RL agents.
-Training various RL agents provides a mechanism for meta-optimization, utilizing the potential of both learning and evolution [25].
-One successful example of PBRL methods is population-based training (PBT) [26], which allows training multiple policies concurrently to enhance the exploration capabilities of the agents in generating diverse behaviors.
-PBT trains a population of agents with different hyperparameters and uses a genetic algorithm to update the population periodically.
-Recently, DexPBT, a decentralized PBRL approach has been proposed to learn dynamic manipulation between two hand-arm systems using parallel simulations [27].
-The authors developed a decentralized implementation to evolve agents in distributed computing environments using on-policy RL, achieving impressive results in dexterous manipulation.
-However, sim-to-real transfer has not been performed, highlighting the complexity of deploying policies on real systems.
-In contrast, this work targets a broader range of real-world tasks including locomotion and manipulation, and transfers the policy onto a real robot without any adaptation phase.
-In addition, the PBRL framework is successfully applied to both off-policy (SAC, DDPG) and on-policy (PPO) RL algorithms, analyzing the implications of critical design choices, i.e., the number of agents and the mutation mechanisms.
+I-A2 Population-Based RL
+Population-based approaches provide a promising solution for exploration and hyperparameter tuning by training a set of policies instead of independent ones.
+By leveraging multiple agents, these methods improve robustness and stabilize training through dynamic hyperparameters adaptation.
+Prior research has demonstrated the effectiveness of such approaches in training deep RL policies, particularly in strategy games and multi-agent interaction [21, 22, 23].
+However, the application of population-based RL (PBRL) methods to robotics remains largely unexplored.
+This limitation arises primarily due to the computational demands and increased training time, which scale linearly with the number of agents when using CPU-based simulators like MuJoCo [24].
+Efficient data collection in such settings typically requires multiple worker machines running separate simulation instances.
+In contrast, Isaac Gym enables large-scale parallel simulation of thousands of robots, significantly accelerating data generation and making it well-suited for training PBRL agents [13].
+A key advantage of training multiple RL agents is its potential for meta-optimization by combining learning and evolutionary mechanisms [25].
+One widely adopted PBRL method is population-based training (PBT) [26], where multiple policies are trained concurrently to enhance exploration and promote diverse behaviors.
+In PBT, an inner loop optimizes policies, while an outer evolutionary loop periodically selects and mutates the best-performing agents.
+More recently, DexPBT [27], a decentralized PBRL approach, has demonstrated impressive results in dexterous manipulation using parallel simulations.
+By distributing evolutionary updates across computing nodes, the authors successfully evolved on-policy RL agents.
+However, no sim-to-real transfer was conducted, underlining the challenges of deploying these policies on physical robots.
+In this paper, we exploit parallel GPU simulation to benchmark PBRL algorithms across 3 RL baselines, including both on-policy (PPO) and off-policy (SAC, DDPG) methods on 4 distinct robotic environments, also evaluating the impact of different hyperparameter mutation mechanisms.
+Furthermore, to increase the significance and extend the applicability of the results, we transfer policies to a real robot, without an additional learning adaptation phase.
 I-A3 Sim-to-Real Transfer
 Despite the calibration efforts to model the physical system accurately, simulation is still a rough approximation.
-The differences between the dynamics of simulated and real systems cause a “reality gap” that makes it unlikely for a simulation-trained policy to successfully transfer to a physical system.
-In literature, researchers have put a significant effort into diminishing this gap: to this aim, most of the approaches leverage domain randomization [4, 6, 16, 17, 28, 29] to expose the policy to a wide range of observation distributions in simulation, thus improving generalization onto a real system.
+The differences between the dynamics of simulated and real systems cause a “reality gap” that makes it unlikely for a simulation-trained policy to successfully transfer to a physical system [28].
+In literature, researchers have put a significant effort into diminishing this gap [29, 30, 31]: to this aim, most of the approaches leverage domain randomization [4, 6, 16, 17, 32, 33] to expose the policy to a wide range of observation distributions in simulation, thus improving generalization onto a real system.
 Nevertheless, naive domain randomization might not be sufficient to completely attenuate the dynamics gap: for instance, [34] employs a specific network to mimic the real actuation system.
 Another technique in this context is policy-level action integrator (PLAI) [10], a simple yet effective algorithm aimed at compensating the sim-to-real dynamic discrepancies with an integral action, which proved to be paramount for a successful transfer.
-In this paper, we employ sim-to-real strategies to deploy a policy on a real system; to the best of the authors’ knowledge, this work represents the first instance of deploying PBRL agents on real hardware.
+In this work, we employ sim-to-real strategies to deploy a policy on a real system; to the best of the authors’ knowledge, this work represents the first instance of deploying PBRL agents on real hardware.
 I-B Contribution
-This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation to dynamically adjust the hyperparameters during training.
-We evaluate the PBRL framework on four complex tasks that require learning essential skills for humanoid robots: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1), available in Isaac Gym [13].
-The results show that better performance is achieved when training a population of agents compared to a single-agent baseline on all tasks.
+This paper investigates a population-based reinforcement learning (PBRL) framework against robotics tasks, allowing to train a population of agents by exploiting GPU-based massively-parallel simulation in order to tune the hyperparameters during training.
+We evaluate the PBRL framework on 4 complex tasks available in Isaac Gym [13]: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Fig.
+1).
+Our results show that training a population of agents yields better performance compared to a single-agent baseline on all tasks.
 The comparison is provided across 3 RL algorithms (PPO, SAC, and DDPG), varying the number of agents in a population, and across different hyperparameter mutation schemes.
-Finally, the PBRL agents are deployed on a real Franka Panda robot for a Franka Nut Pick task, without any policy adaptation phase on the physical system.
+Finally, we deploy the PBRL agents on a real Franka Panda robot for a Franka Nut Pick task, without any policy adaptation phase on the physical system.
 In summary, the main contributions of this work are:
@@
 the number of agents and mutation mechanisms;
-sim-to-real transfer of PBRL policies onto a real Franka Panda robot;
-an open-source codebase to train policies using the PBRL algorithm.
+sim-to-real transfer assessment of PBRL policies onto a real Franka Panda robot;
+an open-source codebase to train policies using the PBRL algorithm: https://github.com/Asad-Shahid/PBRL
 II Methods
 This section describes the core concepts involved in the PBRL framework.
-The overall approach, illustrated in Figure 2, can be viewed as a multi-layered training process consisting of an inner optimization loop with RL and an outer loop of online evolutionary selection with population-based training.
+The overall approach, illustrated in Fig.
+2, can be viewed as a multi-layered training process consisting of an inner optimization loop with RL and an outer loop of online evolutionary selection with population-based training.
 During training, the parameters of the agent’s policy are updated at a higher rate using RL than the hyperparameters defining the RL procedure.
@@
 The RL problem is modeled as a Markov Decision Process (MDP), where an agent interacts with the environment in order to maximize the expected sum of episodic rewards.
-Specifically, an MDP is defined as , where is the set of states, is the set of actions, is the transition dynamics, i.e., , where defines the set of a probability distribution over , is the reward function, and represents the discount factor.
-The goal is formulated as learning a policy, either stochastic, , or deterministic, , where represents a probability distribution over and encapsulates the policy parameters, whose cardinality depends on the selected algorithm and network architecture.
+Specifically, a MDP is defined as , where is the set of states, is the set of actions, is the transition dynamics, where defines a probability distribution over , is the reward function, and represents the discount factor.
+The goal is formulated as learning a policy, either deterministic, , or stochastic, , where represents a probability distribution over , and encapsulates the policy parameters, whose cardinality depends on the selected algorithm and network architecture.
 In this work, the policy is learned using the on-policy method PPO, or either of the off-policy methods SAC or DDPG.
@@
 In DDPG, the common practice involves adding a small noise to the deterministic actions of the policy to enable exploration.
-In this work, the noise is added following a mixed exploration strategy [31], where the general idea is akin to adding a different noise level for each environment when training in a massively parallel regime.
-For the -th environment out of environments, a zero mean and uncorrelated Gaussian noise is given as: , where .
+In this work, the noise is added following a mixed exploration strategy [35], where the general idea is akin to adding a different noise level for each environment when training in a massively-parallel regime.
+For the -th environment out of environments, a zero mean and uncorrelated Gaussian noise is given as: , where , with and being hyperparameters.
 | Hyperparameter | PPO | PBRL-PPO | |||||||
@@
 | Actor variance | 0.5 | 1 | 1 | 0.3 – 1 | 0.3 – 1 | 0.3 – 1 | |||
-| KL threshold | 0.016 | 0.016 | 0.016 | 0.08 – 0.016 | 0.08 – 0.016 | 0.08 – 0.016 | |||
+| KL threshold | 0.016 | 0.016 | 0.016 | 0.008 – 0.016 | 0.008 – 0.016 | 0.008 – 0.016 | |||
 | Entropy loss coefficient | 0.001 | 0.001 | 0 | 0.0001 – 0.001 | 0.0001 – 0.001 | 0.0001 – 0.001 | |||
@@
 II-B Population-Based Training
-In standard RL, the agent aims to learn an optimal policy by interacting with an environment and iteratively updating the policy through some kind of optimization method.
-In contrast, PBRL uses a population of agents , each interacting with the environment independently to collect experience and learn its own policy.
-Using evolutionary selection, the population is periodically evaluated based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
+While a standard RL agent aims to learn an optimal policy by interacting with an environment and iteratively updating the policy through an optimization method, PBRL uses a population composed of agents, each interacting with the environment independently to collect experience and learn its own policy111This setup differs from multi-agent reinforcement learning, where agents interact in a shared environment with a joint or competitive objective.
+Using evolutionary selection, the population is evaluated periodically based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
+While this process does not offer a theoretical guarantee of reaching the global optimum — a cumbersome objective to achieve through naive trial-and-error tuning — evolutionary algorithms have demonstrated a strong empirical tendency to converge to the optimum across a wide range of domains [26, 36, 37].
 In this work, a specific PBRL approach, population-based training (PBT), is employed as an outer optimization loop to enable diverse exploration and dynamically adapt the hyperparameters in high-dimensional continuous control tasks.
-Each agent is characterized by a vector and the set of hyperparameters , where contains the parameters of the policy, and contains the hyperparameters that are optimized during training.
+The -th agent is characterized by the sets and , where contains the parameters of the policy, and contains the hyperparameters that are optimized during training.
 To represent the whole population , we denote with , and the sets of all the parameters, hyperparameters and policies respectively.
@@
 Algorithm 1 provides pseudocode for the PBRL.
-The training process runs in iterations, where all agents are first independently trained by performing updates to the vector .
+Training proceeds iteratively, where all agents are first independently trained by performing updates to the vector .
 After a certain number of policy updates (each agent having been trained for some steps), the agents are evaluated and sorted based on the average return obtained over all of the previous episodes.
 The agents in get replaced by randomly-sampled agents in with mutated hyperparameters, while the rest of the agents in and continue training.
-To generate the mutated hyperparameters, 3 mutation mechanisms are considered (see line 15 of Algorithm 1): (i) random perturbation is applied to the hyperparameters of the parent agent(s) through perturbation factors in Table III; (ii) new hyperparameters are sampled from a prior uniform distribution with bounds specified in Table I and II; (iii) according to the DexPBT mutation scheme [27], hyperparameters are multiplied or divided by a random number sampled from a uniform distribution, i.e., with probability .
-Section III-B4 compares all 3 mutation schemes.
+We consider 3 mutation mechanisms to generate the mutated hyperparameters (see line 15 of Algorithm 1): (i) random perturbation is applied to the hyperparameters of the parent agent(s) through perturbation factors in Table III; (ii) new hyperparameters are sampled from a prior uniform distribution with bounds specified in Table I and II; (iii) according to the DexPBT mutation scheme [27], hyperparameters are multiplied or divided by a random number sampled from a uniform distribution, i.e., with probability .
+Sect.
+III-A4 compares all 3 mutation schemes.
 After beginning the training, evolution is enabled after steps as in [22] to allow for initial exploration and promote population diversity.
 III Experiments
-III-A Environments
-The PBRL framework is evaluated on some of the most challenging benchmark tasks available in Isaac Gym, including Anymal Terrain, Shadow Hand, Humanoid and Franka Nut Pick (Figure 1).
+The PBRL framework is evaluated on some of the most challenging benchmark tasks available in Isaac Gym, including Anymal Terrain, Shadow Hand, Humanoid, and Franka Nut Pick (Fig.
+1).
 The experiments are conducted on a workstation with a single NVIDIA RTX 4090 GPU and of RAM.
 Parallelizing the data collection across the GPU, Isaac Gym’s PhysX engine can simulate thousands of environments using the above hardware.
-III-B Results
+III-A Results
 | Parameter | Value | |
@@
 | Perturbation factor (max.) | 1.2 | 1.2 |
-PPO PBRL-PPO 4 agents
-PBRL-PPO 8 agents PBRL-PPO 16 agents
+Standard RL PBRL 4 agents PBRL 8 agents PBRL 16 agents
 The experiments focus on optimizing the hyperparameters of the RL agents in a population and comparing the results against non-evolutionary baseline agents.
-For each case of baseline agents, 4 experiments are run with different seeds.
-Tables I and II list the hyperparameters for on-policy and off-policy algorithms, including the sampling ranges of those optimized through the PBRL Algorithm 1.
-The initial values for each agent are uniformly sampled from a prior distribution with a given range.
-III-B1 PBRL-PPO
+For each case of baseline agents, 8 experiments are run with different seeds.
+Table I and II provide the hyperparameters for on-policy and off-policy algorithms, listing the sampling ranges of those optimized through the PBRL Algorithm 1.
+For a fair evaluation, we evaluate PBRL on at least 8 agents in total, i.e., for , independent simulations are run, respectively222With a slight abuse of notation, indicates a base RL agent.
+Our benchmarking campaign consists in evaluating the agents’ reward comparing the independent (standard RL) and the evolutionary (PBRL) scenarios, to assess whether agents achieve higher reward or faster convergence when trained in populations.
+III-A1 PBRL-PPO
 For the PPO agents, the tuned hyperparameters are the KL divergence threshold for an adaptive LR, the entropy loss coefficient, and the variance of action selection.
 These parameters are crucial in ensuring sufficient exploration of the environment.
-Figure 3 shows the learning curves for the single-agent PPO baseline and PBRL-PPO for .
-The results demonstrate that PBRL-PPO outperforms PPO on 3 out of 4 tasks, yielding a higher return, with significant improvement seen in Anymal Terrain, which involves traversing increasingly challenging terrain.
-For Franka Nut Pick, PBRL agents achieve comparable performance to the baseline PPO agents.
-This is because, in this relatively straightforward task, randomization alone suffices for a thorough exploration of the state/action space.
-SAC PBRL-SAC 4 agents PBRL-SAC 8 agents
-DDPG PBRL-DDPG 4 agents PBRL-DDPG 8 agents
-III-B2 PBRL-SAC
-In PBRL-SAC, the optimized hyperparameters include the LR of the actor-critic networks and the target entropy factor.
-Entropy is key in SAC agents as the policy is trained to maximize the trade-off between the expected return and exploration.
-Experiments are run with a population size of .
-Due to higher memory needs for replay buffers in off-policy methods, the maximum population size is limited to 8.
-The training performance of SAC and PBRL-SAC is shown in Figure 4.
+Fig.
+3 shows the learning curves for the single-agent PPO baseline and PBRL-PPO for 333Readers should be aware that one can only choose a single agent in at deployment time (e.g.
+the best one), as done in Sect.
+III-C: we compare the agents’ mean reward across to ease the visual comparison.
+The results demonstrate that PBRL-PPO outperforms PPO on 3 out of 4 tasks, yielding a higher return, with significant improvement seen in Anymal Terrain, which involves traversing increasingly challenging terrains.
+For Franka Nut Pick, PBRL agents match baseline PPO performance, as in this relatively straightforward task randomization alone suffices for a thorough exploration of the state/action space.
+III-A2 PBRL-SAC
+The hyperparameters optimized in PBRL-SAC include the LR of the actor-critic networks and the target entropy factor to control the exploration behavior.
+Entropy is a central theme in SAC agents as the policy is trained to maximize the trade-off between the expected return and the entropy.
+The experiments in PBRL-SAC are run with a population size of .
+Off-policy replay buffers require more memory allocation on GPU than on-policy methods.
+Since all agents in PBRL-SAC train with their individual replay buffers, the maximum population size is set to 8 due to the limited GPU memory.
+Fig.
+3 displays SAC and PBRL-SAC training performance.
 PBRL-SAC improves the training performance compared to non-evolutionary SAC on 3 out of 4 tasks, yielding a remarkable improvement on both Shadow Hand and Franka Nut Pick, while comparable results are achieved on Humanoid, probably due to the limited task complexity.
-III-B3 PBRL-DDPG
+III-A3 PBRL-DDPG
 In DDPG, exploration noise is added to the output of the deterministic actor.
@@
 As in PBRL-SAC, the maximum population size in PBRL-DDPG is set to 8 due to the presence of independent replay buffers and GPU memory limitations.
-Figure 5 shows that PBRL-DDPG achieves significantly better training performance than DDPG on all 4 benchmark tasks.
-III-B4 Mutation Comparison
-Figure 6 shows the results using 3 different mutation schemes for PBRL-PPO and PBRL-DDPG.
-As mentioned in Section II-B, the hyperparameters for under-performing agents are generated either by sampling from an original prior distribution, by perturbing the parent’s values through perturbation factors given in Table III, or through the DexPBT mutation scheme presented in [27].
+Fig.
+3 shows that PBRL-DDPG achieves significantly better training performance than DDPG on all 4 benchmark tasks.
+III-A4 Mutation Comparison
+Fig.
+4 shows the results using 3 different mutation schemes for PBRL-PPO and PBRL-DDPG.
+As mentioned in Section II-B, the hyperparameters for under-performing agents are generated either by randomly sampling from their range (see Tables I–II), by perturbing the parent’s values through perturbation factors given in Table III, or through the DexPBT mutation scheme [27].
 In the latter, the hyperparameters have a probability of getting multiplied or divided by a random number sampled from the uniform distribution, .
-The results show that the perturbed agents either exceed or are on par with the performance of other mutation schemes in 6 out of 8 experiments.
-The DexPBT mutation scheme performs better with PBRL-DDPG on Humanoid and Franka Nut Pick tasks, which are less challenging compared to others.
-The combination of two mutation schemes might discover better exploration strategies for a wider range of tasks.
-Analyzing the potential synergies between the two remains a prospect for future investigation.
+The results manifest that the perturbed agents either exceed or are on par with the performance of other mutation schemes in 6 out of 8 experiments, with DexPBT mutation scheme performs better with PBRL-DDPG on Humanoid and Franka Nut Pick tasks, which are less challenging compared to others.
+Generally speaking, the perturbation scheme yields less variance (i.e., a more stable learning) compared to DexPBT.
+III-B Discussion
+Although PBRL agents outperform non-evolutionary agents in nearly all experiments, the influence of population size across different RL algorithms and tasks does not follow a clear or consistent pattern.
+One might intuitively hypothesize that larger, more diverse populations would improve final performance by enabling broader exploration.
+However, our results suggest that increasing the population size does not universally enhance agents training.
+This challenges the common assumption that population-based methods inherently benefit from large populations due to their ability of performing a more thorough exploration of the hyperparameter space [26, 37].
+Instead, the optimal population size appears to be task- and algorithm-dependent, influenced by factors such as task complexity and the nature of agents interaction.
+For simpler tasks, smaller populations may suffice to reach high performance, and additional agents may offer diminishing returns or increase noise during selection.
+While our evolutionary mechanism keeps the degree of exploitation by propagating successful hyperparameters, larger populations may dilute this effect by introducing more diverse, potentially suboptimal, behavior patterns.
+In contrast, tasks with gradually increasing complexity — such as Anymal Terrain, which uses curriculum learning — may benefit more from the exploration enabled by larger populations.
+Additionally, PBRL performance may lag behind non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori.
+This can be noticed on a Humanoid task trained with SAC in Fig.
+3: indeed, baseline policies achieve a higher reward than PBRL-SAC with 4 agents; nevertheless, 8 agents are capable of outperforming the baseline.
+Thus, the benefits provided by PBRL will become more apparent for novel tasks lacking known ideal hyperparameter ranges.
+In this sense, PBRL can be thought of as an exploratory approach to search for unknown optimal configurations of newly designed tasks.
+III-C Sim-to-Real Transfer
+For real-world experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation on a physical robot, deploying the actions with PLAI [10].
+The robot detects the nuts utilizing Mask-RCNN [38], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
 | Parameter | Value |
@@
 | Action scale | 0.0001 |
-III-C Sim-to-Real Transfer
-In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation, executing the actions with PLAI [10].
-The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
-Compared to the original task [4], we applied the following changes to make the simulated environment resemble real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to reflect the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose.
+Compared to the original task [4], we applied the following changes to make the simulated environment as close as possible to the real setup: (i) employing a Task-Space Impedance (TSI) controller [39] instead of an Operational-Space motion Controller (OSC) [40] to comply with the actual low-level controller444The control laws are specified in [4] and in reference works [39, 40]; (ii) randomizing the nut’s initial position to resemble the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose.
 The parameters used in the simulated environment and the real controller are reported in Table IV.
-During experiments, the following policies were deployed, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO.
-With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase.
-As shown in Figure 7, PBRL-PPO with has the highest success rate.
-Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs.
-).
-Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V.
-Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation.
-Unlike the baseline PPO agent, which continued to produce small movements after reaching the target, PBRL-PPO agents remained more stable, leading to a higher success rate.
-This demonstrates that PBRL agents, while achieving similar rewards to a single agent, learn behaviors that exhibit greater robustness to environment variability due to the diversity in agent populations.
-Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
 PPO PBRL-PPO 4 agents
 PBRL-PPO 8 agents PBRL-PPO 16 agents
+During experiments, we deployed the following policies, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO.
+With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase.
+As shown in Fig.
+5, PBRL-PPO with has the highest success rate.
+Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs.
+), also highlighting lower variance compared to both baseline PPO and PBRL-PPO with .
 | Algorithm | Agent | Successful trials | Success rate |
@@
 | PPO | Best | 19/30 | 63.33% |
-III-D Discussion
-While the PBRL agents perform better than the non-evolutionary agents in almost all the experiments, the impact of population size across RL algorithms and tasks shows no consistent pattern.
-One may hypothesize that larger and more diverse populations might lead to a better final performance.
-However, the results in this work indicate that using a larger population size does not necessarily yield substantial benefits for every task.
-This is in contrast to the common belief that population-based methods rely on larger population sizes to effectively explore the hyperparameter space [26, 35].
-The optimal population size, instead, depends on various factors, including task complexity, RL algorithm, and interaction dynamics among agents.
-While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice.
-Larger population sizes seem to perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
-Additionally, the performance of PBRL may be lower than non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori.
-This can be noticed on a Humanoid task trained with SAC in Figure 4: indeed, baseline policies achieve a higher reward than PBRL-SAC with 4 agents; nevertheless, 8 agents are capable of outperforming the baseline.
-Thus, the benefits provided by PBRL will become more apparent for new tasks where ideal hyperparameter ranges are not known in advance.
-In this sense, PBRL can be thought of as an exploratory approach to search for unknown optimal configurations of newly designed tasks.
+Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Fig.
+6), yet with different success rates, as summarized in Table V.
+Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation.
+This suggests that a better exploration across the agents, favored by PBRL, not only leads to higher rewards, but is also one of the aspects that could lead to a successful real-world transfer [31].
+Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
 IV Conclusion
-IV-A Summary
-In this paper, a PBRL framework has been employed to train a population of RL agents by making use of high-throughput GPU-accelerated simulation.
-The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of locomotion and manipulation benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes.
-The results showed the effectiveness of PBRL in improving final performance through online adaptation of hyperparameters.
+In this paper, a PBRL framework has been employed for training a population of RL agents by making use of high-throughput GPU-accelerated simulation.
+The first simulation results of PBRL using on- and off-policy methods are provided on a series of benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes.
+The results showed the effectiveness of PBRL in improving final performance by adapting the hyperparameters online.
 PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning.
-Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms.
-IV-B Future Works
+Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms on challenging robotic tasks.
 Many interesting directions remain for future research.
@@
 This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training.
-Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation or assembly) in the real world with sim-to-real techniques.
-With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable.
+Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation, and assembly) in the real world with sim-to-real techniques.
+With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable, accounting for diverse skills labeling and learning.